Determine the missing amount from each of the separate situations given below. + Equity 37,000 + $ Assets = Liabilities + (a)Required informatione World Of Videos [The following information applies to the questions displayed below.) On October 1, EboRequired information [The following information applies to the questions displayed below.] On October 1, Ebony Ernst organizeRequired information [The following information applies to the questions displayed below. On October 1, Ebony Ernst organizedLita Lopez started Biz Consulting, a new business, and completed the following transactions during its first year of operatio
Answer #1

1)

Assets = Liabilities + Equity
a) (155000+37000)= $192000 = $155000 + $37000
b) $130000 = $59000 + (130000-59000)= $71000
c) $213000 = (213000-62000)= $151000 + $62000

2)

ERNST CONSULTING
Income Statement
For the Month Ended October 31
Revenues
Consulting revenue $14800
Total revenues $14800
Expenses
Rent expense 3750
Salaries expense 7160
Telephone expense 780
Miscellaneous expense 600
Total expenses 12290
Net income $2510
ERNST CONSULTING
Statement of Retained Earnings
For the Month Ended October 31
Retained earnings, Oct 1 $0
Add: Net income 2510
2510
Less: Cash dividends 2180
Retained earnings, Oct 31 $330
ERNST CONSULTING
Balance Sheet
As on October 31
Assets Liabilities
Cash $10710 Accounts payable $8630
Accounts receivable 14800
Office supplies 3450 Equity
Land 46030 Common stock 84200
Office equipment 18170 Retained earnings 330
Total equity 84530
$93160 $93160

3)

Assets Liabilities Equity
Accounts Office Office Accounts Common
Cash + Receivable + Supplies + Equipment + OfficeSuite = Payable + Stock - Dividends + Revenues - Expenses
a. $65000 + + + $36000 + = + $101000 - + -
b. -47000 + + + + 47000 = + - + -
Bal. 18000 + + + 36000 + 47000 = + 101000 - + -
c -5800 + + + 5800 + = + - + -
Bal. 12200 + + + 41800 + 47000 = + 101000 - + -
d + + 3700 + 1900 + = 5600 + - + -
Bal. 12200 + + 3700 + 43700 + 47000 = 5600 + 101000 - + -
e -970 + + + + = + - + - 970
Bal. 11230 + + 3700 + 43700 + 47000 = 5600 + 101000 - + - 970
f + 4300 + + + = + - + 4300 -
Bal. 11230 + 4300 + 3700 + 43700 + 47000 = 5600 + 101000 - + 4300 - 970
g 8100 + + + + = + - + 8100 -
Bal. 19330 + 4300 + 3700 + 43700 + 47000 = 5600 + 101000 - + 12400 - 970
h -1300 + + + + = + - 1300 + -
Bal. 18030 + 4300 + 3700 + 43700 + 47000 = 5600 + 101000 - 1300 + 12400 - 970
i 3300 + -3300 + + + = + - + -
Bal. 21330 + 1000 + 3700 + 43700 + 47000 = 5600 + 101000 - 1300 + 12400 - 970
j -950 + + + + = -950 + - + -
Bal. 20380 + 1000 + 3700 + 43700 + 47000 = 4650 + 101000 - 1300 + 12400 - 970
k -2000 + + + + = + - + - 2000
Bal. $18380 + 1000 + 3700 + 43700 + 47000 = 4650 + 101000 - 1300 + 12400 - 2970